1. Moderate Growth Company paid a dividend last year of $2.10. The expected ROE for next year is 13%. An appropriate required return on the stock is 11%.

If the firm has a plowback ratio of 63%, what should the dividend in the coming year be?
Answer #1

Dividend last year, D0 = $ 2.10

Plow back ratio(Retention ratio) = 63%

Dividend payout ratio = 1- plow back ratio =1- 63% = 37%

37% * EPS = 2.10

EPS last year = 2.1/37% = $ 5.67

Dividend next year = D0*(1+g) where g is the sustainable growth rate

g = ROE * Retention ratio

g = 13% * 0.63 = 8.19%

D1 = 2.10*(1+8.19%) = $ 2.27